OpenStudy (anonymous):

Write the standard equation for the circle center at (0,0) and passing through (squareroot 3/2, 1/2)

OpenStudy (anonymous):

Alright, do you know what standard form is?

OpenStudy (anonymous):

no :[

OpenStudy (anonymous):

The standard form for a circle is \[(x-h)^2+(y-k)^2=r^2\], the coordinates that are given to you are the h and k variables, with the x-coordinate being h and the y-coordinate being k. With that being said, do you think you can set up the left side of the standard form equation with the information you have?

OpenStudy (anonymous):

that would be the "unit circle" the circle with center at the origin, and radius 1

OpenStudy (anonymous):

@satellite73 although you are right, i do believe this is an algebra 2 question

OpenStudy (anonymous):

(x-0)square+(y-0)square=1